JUMBO CONCRETE WALL SUPPORT PIPE TINY PLATFORM PERC PERCH Here is what is known: Jumbo has a weight of 15,000 lbf Tiny has a weight of 5000 lbf. The support tube is fabricated from AISI L-6 tool steel, quenched and tempered at 600 degrees F. (Note: This would be an insanely expensive choice for this tube, AND it would be really hard to weld but the zoo had a big piece of pipe made out of tool steel just laying around so they went ahead and used it.) The support tube has an outer diameter of 16.00 in. and an inner diameter of 15.00 in. Important dimensions are on the next page. FIND THE FACTOR OF SAFETY AGAINST STATIC FAILURE OF THE SUPPORT PIPE
Answer #1

The mechanism of faliure in this problem will be yielding, so we have to consider yield strength of the material to find factor of safety. Generally, yield strength of aisi L6 tool steel is 55114.3 lbf/in2 due to qunched and tempered it will increase upto 130534 lbf/in2.

Since we are talking about static faliure, load on the pipe will be,   

(Load) P = 15000 + 5000 = 20000lbf

(Area of pipe) A = \pi(162-152)/4 = 24.35in2

(Maximum permissible stress) \sigma = P/A = 20000/24.35 = 821.35lbf/in2

Factor of safety = (yield strength)/\sigma = 130534/821.35 = 159